# When to Trust LLMs: Aligning Confidence with Response Quality and Exploring Applications in RAG

### v1

@@ 5.1.2 Baselines
We compare our CONQORD with three baselines on two base models, LLAMA-2 7B and LLAMA-2 13B Touvron et al. (2023). The baselines are as follows:
  • Vanilla elicits verbalized confidence is to directly request them to output a confidence score ranging from 0 to 1.
  • Top-k: Tian et al. Tian et al. (2023) prompts LLMs to generate the top $K$ predictions for a query, each accompanied by an explicit probability that represents the model’s confidence in its prediction.
  • CoT+Agg: Xiong Xiong et al. (2023) leverage the Chain-of-Thought (Wei et al., 2022) prompting strategy. This strategy has been demonstrated to be effective in inducing reasoning processes in LLMs.
@@ 5.2 Evaluation Metric
To evaluate the alignment of the verbalized confidence and response quality, we employ widely used Expected Calibration Error. We also utilize Pearson Correlation coefficient, and Spearman Rank Correlation Coefficient for alignment assessment
@@
Dataset Methods ECE $\downarrow$ Pearson Correlation Spearman Correlation Accuracy $\uparrow$
TruthfulQA Vanilla 0.5887 0.0578 0.0989 0.0616 0.0787 0.3048
Top-k 0.4950 -0.0296 0.3981 0.0055 0.8763 0.4002
CoT+Agg 0.3696 0.0683 0.1272 0.0683 0.1272 0.5100
CONQORD 0.4942 0.0998 0.0043 0.1789 2.7×${\times}\mathbf{10^{-7}}$ 0.3011
NQ Vanilla 0.3892 0.0376 0.4013 0.0376 0.4013 0.5040
Top-k 0.3676 0.0898 0.0448 0.0898 0.0448 0.5100
CoT+Agg 0.3110 0.0778 0.0821 0.0778 0.0821 0.5820
CONQORD 0.2922 0.1005 0.0246 0.116 0.0125 0.4980
Table 2: Alignment performance on base model of LLAMA-2 13B on TruthfulQA and NQ datasets. The symbol $\downarrow$ denotes that lower values are preferable, whereas $\uparrow$ indicates that higher values are more desirable.
@@ 5.2.1 Setup
In our study, we employ the LLAMA 2 architecture for the reward, reference, value, and actor models. The fine-tuning of the reward model utilizes the Helpful & Harmless dataset Bai et al. (2022). Across all models, the AdamW optimizer is the chosen as optimization algorithm. For models with 7B and 13B parameters, we set the KL penalty coefficient, $\beta$, to 0.005, aligning with the parameters used in prior research Touvron et al. (2023). In our primary experiments, we select an $\alpha$ value of 0.4. We apply a weight decay of 0.1 and maintain a constant learning rate of $10^{-6}$. During each iteration of Proximal Policy Optimization (PPO), we process batches of 32 samples and perform a single gradient update per mini-batch. Experiments are conducted on eight 80G A100 GPUs.
@@ 5.3 Confidence Alignment Evaluation
We conduct experiments to demonstrate the effectiveness of our CONCORD on TruthfulQA and NQ datasets. Table 1 and Table 2 illustrate the alignment performance under the base model of LLAMA-2 7B and LLAMA-2 13B. Ideally, a model with lower ECE values, higher Pearson and Spearman coefficients, and lower p-values demonstrates stronger confidence alignment.
We observe that CONQORD generally exhibits strong confidence alignment, as evidenced by the best Expected Calibration Error (ECE) across most of the datasets and model sizes. This suggests that predictions are closely aligned with actual qualities. Our method maintains the base model’s performance while significantly improving calibration, unlike PreApproach, which causes a notable performance drop. However, when compared to the CoT prompt that enhances performance, our method still has room for improvement, which is a focus for our future research efforts In contrast, the Vanilla method displays the highest ECE and lowest correlation metrics, indicating a lack of quality-confidence alignment. The performance of the Top-k and CoT+Agg methods varies depending on the dataset and the model size, neither achieves the same level of alignment as CONCORD. Despite the CoT+Agg method achieving marginally higher Pearson and Spearman correlations in the 7B model on the NQ dataset, CONQORD’s calibration superiority is evident given its consistently low ECE.
@@
5.4 Retrieval Augmentation Experiments
To evaluate the effectiveness of the calibrated CONQORD model in enhancing retrieval augmentation performance, we conducted experiments utilizing the confidence scores generated by the calibrated model. These scores served as a basis for determining whether external retrieval should be employed. Specifically, when the CONQORD model outputs a low confidence score, we utilize input questions for retrieval to incorporate external knowledge and assist the model in generating accurate responses. Conversely, when the confidence score is high, we consider it unnecessary to introduce retrieval at that stage, as the model itself is capable of directly producing high-quality responses.
@@ 5.5 Hyper-Parameter Analysis
As shown in Figure 4, ECE is consistently decreased with the increases of $\alpha$, which validates the effectiveness of $R_{A}$ on aligning the verbalized confidence and response quality. The accuracy is observed to be insensitive to changes in $\alpha$.
@@ 5.6 Case Study
Our case study focusing on TruthfulQA serves as an illustrative example of how our calibrated model, CONQORD, effectively aligns verbalized confidence with the actual quality of responses generated by LLMs. The examples presented in Table 3 reveal that traditional baseline methods frequently overstate confidence, which can mislead users and undermine trust in LLMs. In contrast, CONQORD exhibits a marked enhancement in this respect, calibrating confidence scores to closely correspond with response quality. Where baseline methods might assign unwarranted high confidence to low-quality responses, CONQORD judiciously modulates confidence levels, ensuring that high confidence is indicative of high-quality responses. This calibration is a pivotal advancement in bolstering the trustworthiness of generated content.
@@ 6 Conclusion and Future Work
In this paper, we propose a confidence-quality-order-preserving alignment approach (CONQORD), which marks a significant step forward in the domain of confidence alignment for LLMs. CONQORD is a reinforcement learning strategy with a well-designed dual-component reward function, containing both quality reward and order-preserving alignment reward functions. Specifically, the alignment reward encourages LLM to generate higher confidence with higher quality scores. Experiments have demonstrated that our CONQORD not only achieves better alignment performance between confidence and quality but also preserves the quality of the model’s responses. Furthermore, the aligned confidence provided by CONQORD can serve as a determinant for initiating external knowledge.
@@
This paper proposes a confidence calibration method based on reinforcement learning with human feedback (RLHF) to align the verbalized confidence with actual response quality. However, it is important to acknowledge the limitations of this research. Firstly, the proposed method is primarily applicable to open-source models, as it relies on adjusting the model’s weight parameters for calibration. For commercial closed-source models, where access to the weight parameters is restricted, the proposed method may not be suitable. Additionally, due to practical constraints and experimental costs, this study only conducted experiments on the LLAMA fundation model. Therefore, the generalizability of the proposed method to other models and parameter scales remains unexplored and is left for future work. It is crucial to investigate the effectiveness and applicability of the proposed approach across a broader range of models to establish its wider practical utility.
@@ Appendix A Text Generation Environment
We provide the reinforcement learning environment for text generation in LLM.
Our study concentrates on text generation tasks. For each task, we are provided with a collection of input prompts, denoted as $D=\{x^{n}\}_{n=1}^{N}$. We follow the framework of Ramamurthy et al. (2023); Wu et al. (2023)) to model language generation as a Markov Decision Process (MDP), represented by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{R},P,\gamma,T_{max}\rangle$, where $\mathcal{V}$ is a finite vocabulary set.

### v2

@@ 5.1.2 Baselines
We compare our CONQORD with three baselines on four foundation models, including LLAMA-2 7B, LLAMA-2 13B Touvron et al. (2023), Zephyr 7B Tunstall et al. (2023), and Mistral 7B Jiang et al. (2023). The baseline methods are as follows:
  • Vanilla elicits verbalized confidence to directly request them to output a confidence score ranging from 0 to 1.
  • Top-k: Tian et al. Tian et al. (2023) prompt LLMs to generate the top $K$ predictions for a query, each accompanied by an explicit probability that represents the model’s confidence in its prediction.
  • CoT+Agg: Xiong et al. Xiong et al. (2023) leverage the Chain-of-Thought (Wei et al., 2022) prompting strategy. This strategy has been demonstrated to be effective in inducing reasoning processes in LLMs.
@@ 5.2 Evaluation Metric
To evaluate the alignment of the verbalized confidence and response quality, we employ widely used Expected Calibration Error. We also utilize the Pearson Correlation coefficient and Spearman Rank Correlation Coefficient for alignment assessment:
@@ 5.2.1 Setup
In our study, we employ the foundational model architecture for the reward, reference, value, and actor models. The fine-tuning of the reward model utilizes the Helpful & Harmless dataset Bai et al. (2022). Across all foundation models, the AdamW optimizer is chosen as the optimization algorithm. We set the KL penalty coefficient, $\beta$, to 0.005, aligning with the parameters used in prior research Touvron et al. (2023). In our primary experiments, we select an $\alpha$ value of 0.4. We apply a weight decay of 0.1 and maintain a constant learning rate of $10^{-6}$. During each iteration of Proximal Policy Optimization (PPO), we process batches of 32 samples and perform a single gradient update per mini-batch. Experiments are conducted on eight 80G A100 GPUs.
@@ 5.3 Confidence Alignment Evaluation
We conduct experiments to demonstrate the effectiveness of our CONCORD on TruthfulQA and NQ datasets. Table 1 and Table 2 illustrate the performance of aligning methods under four foundation models (LLAMA-2 7B, Zephyr 7B, Mistral 7B, and LLAMA-2 13B) on the TruthfulQA and NQ datasets. Ideally, a model with lower ECE values, higher Pearson and Spearman coefficients, and lower p-values demonstrates stronger confidence alignment.
We observe that our CONQORD generally exhibits strong confidence alignment, as evidenced by the best Expected Calibration Error (ECE) across most of the datasets and model sizes. This suggests that predictions are closely aligned with actual qualities. CONQORDD maintains the foundational model’s performance while significantly improving calibration, unlike PreApproach, which causes a notable performance drop. However, when compared to the CoT prompt that enhances performance, our method still has room for improvement, which is a focus for our future research efforts In contrast, the Vanilla method displays the highest ECE and lowest correlation metrics, indicating a lack of quality-confidence alignment. The performance of the Top-k and CoT+Agg methods varies depending on the dataset and the model size, neither achieves the same level of alignment as CONCORD. Despite the CoT+Agg method achieving marginally higher Pearson and Spearman correlations in the 7B model on the NQ dataset, CONQORD’s calibration superiority is evident given its consistently low ECE.
@@
5.4 Adaptive Retrieval Experiments
To evaluate the effectiveness of the calibrated CONQORD model in enhancing retrieval augmentation performance, we conduct experiments utilizing the confidence scores generated by the calibrated model. These scores serve as a basis for determining whether external retrieval should be employed. Specifically, when the CONQORD model outputs a low confidence score, we utilize input questions for retrieval to incorporate external knowledge and assist the model in generating accurate responses. Conversely, when the confidence score is high, we consider it unnecessary to introduce retrieval at that stage, as the model itself is capable of directly producing high-quality responses.
@@ 5.5 Hyper-Parameter Analysis
As shown in Figure 4, ECE is decreased with the increases of $\alpha$, which validates the effectiveness of alignment reward $R_{A}$ on aligning the verbalized confidence and response quality. The accuracy is observed to be insensitive to changes in $\alpha$.
@@ 5.6 Case Study
Our case study focusing on TruthfulQA serves as an illustrative example of how our calibrated model, CONQORD, effectively aligns verbalized confidence with the actual quality of responses generated by LLMs. The examples presented in Table 3 reveal that traditional baseline methods frequently overstate confidence, which can mislead users and undermine trust in LLMs. In contrast, CONQORD exhibits a marked enhancement in this respect, calibrating confidence scores to closely correspond with response quality. Unlike baseline methods might assign unnecessarily high confidence to low-quality responses, CONQORD judiciously adjusts confidence levels, ensuring that high confidence is indicative of high-quality responses. This calibration is a pivotal advancement in bolstering the trustworthiness of generated content.
@@ 6 Conclusion and Future Work
In this paper, we propose a confidence-quality-order-preserving alignment approach (CONQORD), which marks a significant step forward in the domain of confidence alignment for LLMs. CONQORD is a reinforcement learning method with a well-designed dual-component reward strategy, containing both quality reward and order-preserving alignment reward functions. Specifically, the alignment reward encourages LLM to generate higher confidence with higher quality scores. Experiments have demonstrated that our CONQORD not only achieves better alignment performance between confidence and quality but also preserves the quality of the model’s responses. Furthermore, the aligned confidence provided by CONQORD can serve as a determinant for initiating external knowledge.
@@
This paper proposes a confidence calibration method based on reinforcement learning with human feedback (RLHF) to align the verbalized confidence with actual response quality. However, it is important to acknowledge the limitations of this research. Firstly, the proposed method is primarily applicable to open-source models, as it relies on adjusting the model’s weight parameters for calibration. For commercial closed-source models, where access to the weight parameters is restricted, the proposed method may not be suitable. Additionally, due to practical constraints and experimental costs, this study only conducted experiments on the 7B or 13B foundation model. Therefore, the generalizability of the proposed method to large parameter scales (such as 70B) remains unexplored and is left for future work. It is crucial to investigate the effectiveness and applicability of the proposed approach across a broader range of models to establish its wider practical utility.
Acknowledgement
This work is funded by the National Key R&\&&D Program of China (2022YFB3103700, 2022YFB3103701), the Strategic Priority Research Program of the Chinese Academy of Sciences under Grant No. XDB0680101.
@@ Appendix A Text Generation Environment
We provide a reinforcement learning environment for text generation in LLM.
Our study concentrates on text-generation tasks. For each task, we are provided with a collection of input prompts, denoted as $D=\{x^{n}\}_{n=1}^{N}$. We follow the framework of Ramamurthy et al. (2023); Wu et al. (2023)) to model language generation as a Markov Decision Process (MDP), represented by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{R},P,\gamma,T_{max}\rangle$, where $\mathcal{V}$ is a finite vocabulary set.